# Optimal Decision-Making Based on Prediction Sets

### v2

@@ 1 Introduction
Predictions are rarely an end in themselves: they are used to select actions. In medicine, a model’s diagnosis informs treatment; in robotics, perception informs control. In such settings, the downstream objective is not “accuracy” per se but low decision loss. The classical statistical decision-theoretic prescription is clear: if the conditional law of the outcomes $Y$ given the features $x$, $P(Y\mid X=x)$, were known, one would choose the Bayes action minimizing expected loss $\mathbb{E}[\ell(a,Y)\mid X=x]$ (Wald, 1945, 1949; Lehmann and Casella, 1998). In modern ML pipelines, however, the conditional law is unknown and predictive models are imperfect. This has motivated a great deal of work on optimal decision making under uncertainty, see e.g., Keith and Ahner (2021); Elmachtoub and Grigas (2022); Zhao et al. (2021), etc. However, many of these works still assume that one can obtain “correct” probability predictions in some weak sense, e.g., consistent or calibrated predictions.
@@
A finite-sample algorithm with distribution-free coverage. To translate the oracle characterization into a practical procedure, we introduce Risk-Optimal Conformal Prediction (ROCP). ROCP uses any black-box probabilistic model to construct estimates of the population quantities in our oracle decision rule, and then uses a held-out calibration set to ensure coverage via conformal prediction (Algorithm 1). This guarantees finite-sample marginal coverage under exchangeability (Theorem 4.1), while targeting the decision-theoretic optimum as the model improves. Empirically, ROCP consistently reduces worst-case risks and critical mistake rates relative to risk-averse max–min baselines and best-response approaches, with the largest gains arising precisely in regimes where the $\alpha$ fraction of out-of-set mass can induce catastrophic losses.
ML model Prediction set design $C^{*}(x)$ Action policy design $a^{*}(x)$ Optimal action policy(Thm. 2.2)Optimal prediction sets(Thm. 3.3, Prop. 4.1)Uncertainty quantificationDecision makingPredictionPre-trainedRisk-Optimal Conformal Prediction (ROCP)
Figure 1: The pipeline of Risk-Optimal Conformal Prediction (ROCP).
Together, these ideas provide an approach to connect prediction sets and actions, as illustrated in Figure 1.
@@ 2 Optimal action selection from prediction sets
2.1 Problem Formulation
In this section, we formalize the goal of minimizing the expected loss, leveraging a prediction set. Let $\mathcal{X}$ and $\mathcal{Y}$ be the feature and outcome spaces, and let $\mathcal{A}$ be the action space. For instance, the actions could be the outcomes: $\mathcal{A}=\mathcal{Y}$.
@@ 2.2 Optimal Policies
If for every $x\in\mathcal{X}$, the function $a\mapsto L_{C(x)}(a;\alpha)$ attains its minimum  [^5] If $\mathcal{A}$ is compact and $a\mapsto\ell(a,y)$ is lower semicontinuous for each $y$, then the function $a\mapsto L_{S}(a;\alpha)$ attains its minimum., then optimal policies $\pi^{\star}$ to the problem (1) have the form  [^6] Existence of measurable selection is discussed in Appendix B; see Remark B.1 after the proof of Theorem 2.2.
@@
3 The Optimal Prediction Sets
In this section, we characterize the optimal prediction sets for the problem (6) by first studying the pointwise-across-$x$ problem and introducing a reparametrization (see (9)) which we then solve using techniques from duality theory.
Starting from this section, we drop the subscript $P$ on $\mathbb{P}_{P}$ and $\mathbb{E}_{P}$ for simplicity. Our analysis begins with a pointwise problem: fix a feature value $x\in\mathcal{X}$ and a target certified conditional coverage level $t\in(0,1]$. We will design a set $C(x)\subseteq\mathcal{Y}$ subject to the constraint $\Pr\{Y\in C(x)\mid X=x\}\geqslant t$, and evaluate decisions under the corresponding miscoverage budget 1t1-t via the robust risk $R(C(x),1-t)=\min_{a\in\mathcal{A}}L_{C(x)}(a;1-t)$.
@@ 3 The Optimal Prediction Sets
Also, consider the loss sublevel set for some $\theta$, namely $S_{\theta}(a)\ :=\ \{y\in\mathcal{Y}:\ \ell(a,y)\leqslant\theta\}$. Intuitively, $S_{Q_{t}^{x}(a)}(a)$ are the lowest-loss labels under action $a$, with $x$-conditional coverage of at least $t$. This is the natural candidate feasible set with small worst-case in-set loss $\sup_{y\in C(x)}\ell(a,y)$. Moreover, under the constraint $\mathbb{P}(Y\in C(x)\mid X=x)\geqslant t$, an adversary in the definition of $L_{C(x)}(a;1-t)$ may place probability $t$ on a worst in-set label (with loss $Q_{t}^{x}(a)$ when $C(x)=S_{Q_{t}^{x}(a)}(a)$) and the remaining probability 1t1-t on a label achieving the maximal loss $M(a)$, leading to the pointwise objective $t\,Q_{t}^{x}(a)\ +\ (1-t)\,M(a)$. Then the corresponding pointwise optimal action is
@@
The associated threshold below which the loss values are included is $\theta(x,t):=Q_{t}^{x}(a(x,t))$; and the resulting optimal set $C(x,t)$ is
@@
The following proposition summarizes this formally.
@@
Fix an instance $x\in\mathcal{X}$ and a coverage value $t\in(0,1]$. Assume the minimum in (7) exists, [^7] This holds, e.g., if $\mathcal{A}$ is compact and $a\mapsto Q_{t}^{x}(a)$ and $a\mapsto M(a)$ are lower semicontinuous. then among all the sets $C\subseteq\mathcal{Y}$ that have coverage at least $t$, i.e. $\operatorname{Pr}(Y\in C\mid X=x)\geq t$, the set $C(x,t)$ from (8) achieves the smallest risk R(C,1t)R(C,1-t) with
@@
Given the above proposition, the optimal set for (6) can be obtained based on the following re-parametrization in terms of the coverage probabilities that we assign to each $x\in\mathcal{X}$. In order to satisfy the marginal constraint, we require $\mathbb{E}_{X}[t(X)]\geq 1-\alpha$. From the above proposition, we know that the pointwise optimal set is $C(x,t)$ with $R(C(x,t),1-t)\ =\ t\,\theta(x,t)\ +\ (1-t)\,M\big(a(x,t)\big)$. Hence, to find the optimal prediction sets, we will need to find the assignment $t(x)$ which optimally distributes the $1-\alpha$ probability over the feature space $\mathcal{X}$, such that the expected risk is optimized. Formally, the problem has the following equivalent reformulation:
@@
and set $V_{x}(0):=\min_{a\in\mathcal{A}}M(a)$ as in Remark 3.1. Once the optimal solution $t^{\star}(x)$ to the above re-parametrization is found, then the optimal policy/actions, denoted by $a^{\star}(x)=a\left(x,t^{\star}(x)\right)$, are derived according to (7), and the optimal prediction set is given by:
@@
Using tools from duality theory, we can show that the above optimization problem admits a solution with a simple ”one-dimensional” structure in terms of scalar parameter $\beta\in\mathbb{R}_{+}$ and an selector function $g:\mathcal{X}\times\mathbb{R}_{+}\rightarrow[0,1]$ defined as [^8] For simplicity in the main text, we assume that for every $\beta\geqslant 0$ the argmin in (12) is a singleton for $P_{X}$-a.e. $x$. The proof of Theorem 3.3 in Appendix B gives the general argument that allows pointwise ties.
$g(x,\beta)\ =\ \arg\min_{t\in[0,1]}\ \big\{\,V_{x}(t)\ -\ \beta\,t\,\big\},$ (12)
@@
Fix $\alpha\in(0,1)$. Assume $P_{X}$ is non-atomic, that $(x,t)\mapsto V_{x}(t)$ in (10) is a normal integrand, [^9] See Appendix B for the definition of normal integrands and its equivalent characterization. and that, for each $x$, the minimum in the definition of $V_{x}(t)$ exists for every $t\in[0,1]$. Then there exists $\beta^{\star}\geqslant 0$ such that
$t^{\star}(x)\ :=\ g(x,\beta^{\star})$
solves the population problem (9). Consequently, the optimal prediction sets are obtained using $t^{\star}(x)$ from (11), with actions $a^{\star}(x)=a\big(x,t^{\star}(x)\big)$ as in (7).
Moreover, $\beta^{\star}$ is characterized by the scalar equation
$\mathbb{E}_{X}\big[g(X,\beta)\big]\ =\ 1-\alpha,$ (13)
in the sense that any optimal $\beta^{\star}>0$ must satisfy (13), and conversely, any $\beta\geqslant 0$ satisfying (13) yields an optimal $t^{\star}(x)=g(x,\beta)$. If (13) has no solution with $\beta>0$ but $\mathbb{E}[g(X,0)]\geqslant 1-\alpha$, then $\beta^{\star}=0$.
@@ 4.1 Constructing estimators
We assume access to a set of calibration data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and a predictive model $f:\mathcal{X}\mapsto\Delta_{\mathcal{Y}}$. The output of the predictive model $f$ is a probability distribution over $\mathcal{Y}$ that estimates the true conditional distribution $P(y|x)$. We use the notation $f_{x}(y)$ to denote the probability assigned to label $y$ for the input $x$. Next, we develop an algorithm leveraging this model to take good actions based on a prediction set.
Following the oracle construction from Section 3, we first estimate $Q_{t}^{x}(a)$ using $f_{x}$:
$\hat{Q}_{t}^{x}(a):=\inf\big\{\theta\in\mathbb{R}:\mathbb{P}_{Y\sim f_{x}}(\ell(a,Y)\leqslant\theta)\geqslant t\big\},$
then estimate
$\widehat{a}(x,t)\in\arg\min_{a\in\mathcal{A}}\Big\{t\widehat{Q}_{t}^{x}(a)+(1-t)M(a)\Big\},$
and set $\widehat{\theta}(x,t):=\widehat{Q}_{t}^{x}\big(\widehat{a}(x,t)\big)$. Finally, define the empirical counterpart of $V_{x}$,
$\widehat{V}_{x}(t):=\min_{a\in\mathcal{A}}\Big\{t\,\widehat{Q}_{t}^{x}(a)+(1-t)\,M(a)\Big\},$
and the empirical dual selector:
$\widehat{g}(x,\beta)\in\arg\min_{t\in[0,1]}\ \big\{\widehat{V}_{x}(t)-\beta t\big\}.$
From Theorem 3.3, we know that the optimal prediction sets admit a ”one-dimensional” structure in terms of the scalar parameter $\beta\geqslant 0$, and the optimal coverage assignment is derived using the function $g$. Hence, to simplify notation, we define the $\beta$–parametrized quantities:
$\widehat{\theta}(x,\beta)\ :=\ \widehat{\theta}\!\big(x,\widehat{g}(x,\beta)\big),\quad\widehat{a}(x,\beta)\ :=\ \widehat{a}\!\big(x,\widehat{g}(x,\beta)\big).$
Following (11), the plug-in estimate of the optimal set takes the form $\widehat{C}(x;\beta):=\widehat{C}_{0}\big(x;\widehat{g}(x,\beta)\big)$, where
$\widehat{C}_{0}(x;t):=\Big\{y\in\mathcal{Y}:\ \ell\big(\widehat{a}(x,t),y\big)\leqslant\widehat{\theta}(x,t)\Big\}.$ (14)
@@ 4.2 Risk-optimal conformal prediction
1:Miscoverage level $\alpha\in(0,1)$, calibration samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$, test covariate $X_{\rm test}$.
@@
Proposition 4.1 (Distribution-free marginal coverage).
Assume the calibration pairs $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and the test pair $(X_{\rm test},Y_{\rm test})$ are exchangeable. Then the ROCP set satisfies
$\Pr\!\big\{\,Y_{\rm test}\in C_{\rm ROCP}(X_{\rm test})\,\big\}\ \geqslant\ 1-\alpha,$
where the probability is over the joint randomness of the calibration and test data.
@@ 5 Experiments
    Average Realized Worst-Case Risk: the mean of the worst-case robust risk $\frac{1}{n_{\rm test}}\sum_{i=1}^{n_{\rm test}}L_{C(x_{i})}(a_{i};\alpha),$ where $L_{S}(a;\alpha)$ is defined in Lemma 2.1 and $a_{i}$ is the chosen action.
@@ Appendix B Proofs
We prove a general statement of Theorem 3.3 where we do not assume pointwise uniqueness of the minimizer in (12).
Theorem B.3.
Assume $P_{X}$ is non-atomic, that $(x,t)\mapsto V_{x}(t)$ in (10) is a normal integrand, and that, for each $x$, the minimum in the definition of $V_{x}(t)$ exists for every $t\in[0,1]$. Then there exists $\beta^{*}\geqslant 0$ and a measurable $t^{\star}:\mathcal{X}\to[0,1]$ such that
$t^{\star}(x)\in\Gamma_{\beta^{\star}}(x)\quad\text{$P_{X}$-a.e.\ $x$},$
and $t^{*}$ solves the population problem (9). Consequently, the optimal prediction sets are obtained using $t^{\star}(x)$ from (11), with actions $a^{\star}(x)=a\big(x,t^{\star}(x)\big)$ as in (7). Moreover, one can always choose $t^{*}$ of the form
Proof of Theorem B.3.
@@
Proof of Theorem 3.3.
In the main text we assume that for every $\beta\geqslant 0$, the argmin in (12) is a singleton for $P_{X}$-a.e. $x$. Thus, for each fixed $\beta$, $\Gamma_{\beta}(x)=\{g(x,\beta)\}$ for $P_{X}$-a.e. $x$, and
$g^{-}(x,\beta)=g^{+}(x,\beta)=g(x,\beta),\text{ a.s. }$
Then our general form (16) reduces to $t^{\star}(x)=g(x,\beta^{\star})$.
[…]
$\mathbb{E}\big[g(X,\beta^{\star})\big]\ =\ 1-\alpha,$
which is exactly (13).
Conversely, suppose some $\beta\geqslant 0$ solves (13), and define $t_{\beta}(x):=g(x,\beta)$. Then $t_{\beta}(x)\in\Gamma_{\beta}(x)$ for $P_{X}$-a.e. $x$ and $\mathbb{E}[t_{\beta}(X)]=1-\alpha$. From the proof of Theorem B.3, the dual function $\phi$ is concave and its one-sided derivatives satisfy
$\phi^{\prime}_{+}(\beta)=(1-\alpha)-\mathbb{E}\left[g^{+}(X,\beta)\right],\qquad\phi^{\prime}_{-}(\beta)=(1-\alpha)-\mathbb{E}\left[g^{-}(X,\beta)\right].$
Under the singleton assumption, $g^{+}(X,\beta)=g^{-}(X,\beta)=t_{\beta}(X)$ a.s., hence
$\phi^{\prime}_{+}(\beta)=\phi^{\prime}_{-}(\beta)=(1-\alpha)-\mathbb{E}\left[t_{\beta}(X)\right]=0.$
If $\beta>0$, this implies that $\beta$ is a maximizer of $\phi$. If $\beta=0$, concavity of $\phi$ and $\phi^{\prime}_{+}(0)=0$ imply $\phi(\beta^{\prime})\leqslant\phi(0)$ for all $\beta^{\prime}\geqslant 0$, so 0 is also a maximizer. By strong duality (Lemma B.5), $\operatorname{VAL}(\alpha)=\sup_{\beta^{\prime}\geqslant 0}\phi(\beta^{\prime})=\phi(\beta)=\mathbb{E}\!\left[V_{X}\big(t_{\beta}(X)\big)\right]$, so $t_{\beta}(\cdot)=g(\cdot,\beta)$ is primal optimal. Therefore, if (13) has multiple solutions, any $\beta^{\star}\geqslant 0$ satisfying it yields an optimal $t^{\star}(x)=g(x,\beta^{\star})$.
Finally, if (13) has no solution with $\beta>0$ but $\mathbb{E}[g(X,0)]\geqslant 1-\alpha$, then necessarily $\mathbb{E}[g(X,0)]>1-\alpha$ (otherwise $\beta=0$ would solve (13)). In this case there cannot exist a dual maximizer with $\beta^{\star}>0$ (since any such maximizer would satisfy (13)), hence any dual maximizer must satisfy $\beta^{\star}=0$, and the primal average-coverage constraint is slack. ∎
Proof of Proposition 4.1. Let $Z_{i}=(X_{i},Y_{i})$ for $i=1,\dots,n$ and $Z_{n+1}=(X_{\rm test},Y_{\rm test})$. For any $\beta\geqslant 0$ define the empirical coverage fraction
$\widehat{p}(\beta):=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbf{1}\{Y_{i}\in\widehat{C}(X_{i};\beta)\}.$
By definition of Algorithm 1, when we plug in the candidate label $y=Y_{\rm test}$ we obtain
$\widehat{\beta}_{Y_{\rm test}}=\inf\{\beta\geqslant 0:\ \widehat{p}(\beta)\geqslant 1-\alpha\},$
where we select the smallest feasible $\beta$. Since $\widehat{p}(\beta)$ is symmetric in the n+1n{+}1 pairs $\{Z_{i}\}_{i=1}^{n+1}$, the random variable $\widehat{\beta}_{Y_{\rm test}}$ is also a symmetric measurable function of $\{Z_{i}\}_{i=1}^{n+1}$.
Now define $I_{i}:=\mathbf{1}\{Y_{i}\in\widehat{C}(X_{i};\widehat{\beta}_{Y_{\rm test}})\}$ for $i=1,\dots,n+1$. Because the full collection $\{Z_{i}\}_{i=1}^{n+1}$ is exchangeable and the map from $\{Z_{i}\}$ to $(I_{1},\dots,I_{n+1})$ is permutation-equivariant, $(I_{1},\dots,I_{n+1})$ is an exchangeable $\{0,1\}^{n+1}$-valued vector. Therefore, $\mathbb{E}[I_{n+1}]=\mathbb{E}[(n{+}1)^{-1}\sum_{i=1}^{n+1}I_{i}]$. Finally, by the defining constraint for $\widehat{\beta}_{Y_{\rm test}}$ we have $\widehat{p}(\widehat{\beta}_{Y_{\rm test}})\geqslant 1-\alpha$, i.e. $(n{+}1)^{-1}\sum_{i=1}^{n+1}I_{i}\geqslant 1-\alpha$, hence
$\Pr\{Y_{\rm test}\in C_{\rm ROCP}(X_{\rm test})\}=\mathbb{E}[I_{n+1}]=\mathbb{E}\!\left[\frac{1}{n+1}\sum_{i=1}^{n+1}I_{i}\right]\geqslant 1-\alpha.$

### v3

@@ 1 Introduction
When making predictions, we often need to take into account how they affect downstream decisions. The classical statistical decision-theoretic prescription is clear: if the conditional law of the outcomes $Y$ given the features $x$, $P(Y\mid X=x)$, were known, one would choose the Bayes action minimizing expected loss $\mathbb{E}[\ell(a,Y)\mid X=x]$ (Wald, 1945, 1949; Lehmann and Casella, 1998). In modern ML pipelines, however, the conditional law is unknown and predictive models are imperfect. This has motivated a great deal of work on optimal decision making under uncertainty, see e.g., Keith and Ahner (2021); Elmachtoub and Grigas (2022); Zhao et al. (2021), etc. However, many of these works still assume that one can obtain “correct” probability predictions in some weak sense, e.g., consistent or calibrated predictions.
@@
A finite-sample algorithm with distribution-free coverage. To translate the oracle characterization into a practical procedure, we introduce Risk-Optimal Conformal Prediction (ROCP). ROCP uses any black-box probabilistic model to construct estimates of the population quantities in our oracle decision rule, and then uses a held-out calibration set to ensure coverage via conformal prediction (Algorithm 1). This guarantees finite-sample marginal coverage under exchangeability, while targeting the decision-theoretic optimum as the model improves. Empirically, ROCP consistently reduces worst-case risks and critical mistake rates relative to risk-averse max–min baselines and best-response approaches, with the largest gains arising precisely in regimes where the $\alpha$ fraction of out-of-set mass can induce catastrophic losses.
@@ 2 Optimal action selection from prediction sets
2.1 Setting
We study how to minimize an expected loss, leveraging a prediction set. Let $\mathcal{X}$ and $\mathcal{Y}$ be the feature and outcome spaces, and let $\mathcal{A}$ be the action space. For instance, the actions could be the outcomes: $\mathcal{A}=\mathcal{Y}$.
@@ 2.2 Optimal Policies
If for every $x\in\mathcal{X}$, the function $a\mapsto L_{C(x)}(a;\alpha)$ attains its minimum, [^5] If $\mathcal{A}$ is compact and $a\mapsto\ell(a,y)$ is lower semicontinuous for each $y$, then the function $a\mapsto L_{S}(a;\alpha)$ attains its minimum. then optimal policies $\pi^{\star}$ to the problem (1) have the form [^6] Existence of measurable selection is discussed in Appendix B; see Remark B.1 after the proof of Theorem 2.2.
@@
3 Oracle Optimality
Here, we study the oracle optimal sets from (6). We start by following the approach from Kiyani et al. (2025), and then make the necessary changes to our setting.
Starting from this section, we drop the subscript $P$ on $\mathbb{P}_{P}$ and $\mathbb{E}_{P}$ for simplicity. Our analysis begins with a pointwise problem: fix a feature value $x\in\mathcal{X}$ and a target certified conditional coverage level $t\in(0,1]$. As in Kiyani et al. (2025), we will design a set $C(x)\subseteq\mathcal{Y}$ subject to the constraint $\Pr\{Y\in C(x)\mid X=x\}\geqslant t$, and evaluate decisions under the corresponding miscoverage budget 1t1-t via the robust risk $R(C(x),1-t)=\min_{a\in\mathcal{A}}L_{C(x)}(a;1-t)$.
@@ 3 Oracle Optimality
Also, consider the loss sublevel set for some $\theta$, namely $S_{\theta}(a):=\{y\in\mathcal{Y}:\ \ell(a,y)\leqslant\theta\}$. Intuitively, $S_{Q_{t}^{x}(a)}(a)$ are the lowest-loss labels under action $a$, with $x$-conditional coverage of at least $t$. This is the natural candidate feasible set with small worst-case in-set loss $\sup_{y\in C(x)}\ell(a,y)$.
Moreover, under the constraint $\mathbb{P}(Y\in C(x)\mid X=x)\geqslant t$, an adversary in the definition of $L_{C(x)}(a;1-t)$ may place probability $t$ on a worst in-set label (with loss $Q_{t}^{x}(a)$ when $C(x)=S_{Q_{t}^{x}(a)}(a)$) and the remaining probability 1t1-t on a label achieving the maximal loss $M(a)$, leading to the pointwise objective $t\,Q_{t}^{x}(a)\ +\ (1-t)\,M(a)$. Then the corresponding pointwise optimal action is
@@
This is analogous to the optimal action $a$ from Kiyani et al. (2025) for the quantile setting. The associated threshold, below which the loss values are included, is $\theta(x,t):=Q_{t}^{x}(a(x,t))$; and the resulting optimal set $C(x,t)$ is
@@
The following proposition summarizes this formally. It is an analogue of Proposition 3.1 in Kiyani et al. (2025), with the difference that our formulation for expected risk necessitates explicitly accounting for the (1t)(1-t) probability mass falling outside the set, yielding the additional (1t)M(a)(1-t)M(a) term absent from their formulation.
@@
If the minimum in (7) exists, [^7] This holds, e.g., if $\mathcal{A}$ is compact and $a\mapsto Q_{t}^{x}(a)$ and $a\mapsto M(a)$ are lower semicontinuous. then, subject to $\operatorname{Pr}(Y\in C\mid X=x)\geq t$, the set $C(x,t)$ from (8) achieves the smallest risk R(C,1t)R(C,1-t) with
@@
Proposition 3.1 allows us to reparametrize the problem (6) in terms of the pointwise coverage $t:x\mapsto t(x)$, as in Kiyani et al. (2025). Formally, the problem (6) has the following equivalent reparametrization:
@@
and set $V_{x}(0):=\min_{a\in\mathcal{A}}M(a)$ as in Remark 3.1. Letting $t^{\star}$ be the optimum, the optimal actions are $a^{\star}(x)=a\left(x,t^{\star}(x)\right)$, and the optimal prediction set is:
@@
To solve this problem, we adopt a duality-based approach in the spirit of Kiyani et al. (2025). However, their proof strategy does not transfer directly to our setting. They reduce the problem to an infinite-dimensional linear program and leverage LP duality theory, which does not work for our problem. Instead, we introduce a different technical approach: we work in the normal-integrand framework and invoke Fenchel-Rockafellar duality through a randomized-kernel relaxation, followed by derandomization to recover a deterministic optimizer.
For $\beta\geqslant 0$, we define the dual function
and define the (set-valued) argmin correspondence $\Gamma_{\beta}(x):=\arg\min_{u\in[0,1]}\bigl\{V_{x}(u)-\beta u\bigr\},$ and its extremal selectors
The following theorem characterizes the optimal $t^{*}$ and hence the optimal action and prediction set.
@@
Assume $P_{X}$ is non-atomic, that $(x,t)\mapsto V_{x}(t)$ in (10) is a normal integrand [^8] See Appendix B for the definition of normal integrands and its equivalent characterization., and that, for each $x$, the minimum in the definition of $V_{x}(t)$ exists for every $t\in[0,1]$. Then there exists $\beta^{*}\geqslant 0$ and a measurable $t^{\star}:\mathcal{X}\to[0,1]$ such that
$t^{\star}(x)\in\Gamma_{\beta^{\star}}(x)\quad\text{$P_{X}$-a.e.\ $x$},$
and $t^{*}$ solves the population problem (9). Substituting this optimal coverage assignment $t^{\star}(x)$ into (11) yields the optimal prediction sets, with actions $a^{\star}(x)=a\big(x,t^{\star}(x)\big)$ as in (7). Moreover, one can always choose $t^{*}$ of the form
@@ 4.1 Constructing estimators
Given calibration data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and a probabilistic predictor $f:\mathcal{X}\mapsto\Delta_{\mathcal{Y}}$, where we consider $f_{x}(y)$ to estimate the true conditional distribution $P(y|x)$. We adopt the approach of estimating the population-level quantities with plug-in estimators and then conformalizing the final result to guarantee coverage. Such approaches have been widely used in conformal prediction, see e.g., Sadinle et al. (2019); Romano et al. (2020); Kiyani et al. (2025); Wang et al. (2025), etc.
Following the oracle construction from Section 3, as in Sadinle et al. (2019); Romano et al. (2020); Kiyani et al. (2025); Wang et al. (2025), we first estimate $Q_{t}^{x}(a)$ using a standard plug-in principle leveraging $f_{x}$: $\hat{Q}_{t}^{x}(a):=\inf\big\{\theta\in\mathbb{R}:\mathbb{P}_{Y\sim f_{x}}(\ell(a,Y)\leqslant\theta)\geqslant t\big\},$ then estimate $\widehat{a}(x,t)\in\arg\min_{a\in\mathcal{A}}\Big\{t\widehat{Q}_{t}^{x}(a)+(1-t)M(a)\Big\},$ and set $\widehat{\theta}(x,t):=\widehat{Q}_{t}^{x}\big(\widehat{a}(x,t)\big)$. Finally, define the plug-in estimate of $V_{x}$:
$\widehat{V}_{x}(t):=\min_{a\in\mathcal{A}}\Big\{t\,\widehat{Q}_{t}^{x}(a)+(1-t)\,M(a)\Big\}.$
and the empirical dual selector: $\widehat{g}(x,\beta)\in\arg\min_{t\in[0,1]}\ \big\{\widehat{V}_{x}(t)-\beta t\big\}.$ Following Theorem 3.3 and as in Kiyani et al. (2025), we define the $\beta$–parametrized quantities: $\widehat{\theta}(x,\beta):=\widehat{\theta}\!\big(x,\widehat{g}(x,\beta)\big)$, $\widehat{a}(x,\beta):=\widehat{a}\!\big(x,\widehat{g}(x,\beta)\big)$. Following (11), the plug-in estimate of the optimal set is $\widehat{C}(x;\beta):=\widehat{C}_{0}\big(x;\widehat{g}(x,\beta)\big)$, where $\widehat{C}_{0}(x;t):=\Big\{y\in\mathcal{Y}:\ \ell\big(\widehat{a}(x,t),y\big)\leqslant\widehat{\theta}(x,t)\Big\}$.
@@ 4.2 Risk-optimal conformal prediction
1:Calibration samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$, test covariate $X_{\rm test}$.
@@
As in Gibbs et al. (2025); Kiyani et al. (2025), it follows that if the test datapoint $(X_{\rm test},Y_{\rm test})$ and the calibration data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ are exchangeable, the ROCP set satisfies the standard marginal coverage guarantee (Vovk et al., 1999, 2005) $\Pr\!\big\{Y_{\rm test}\in C_{\rm ROCP}(X_{\rm test})\,\big\}\ \geqslant\ 1-\alpha,$
@@ 5 Experiments
    Average Realized Worst-Case Risk: the mean worst-case robust risk $\frac{1}{n_{\rm test}}\sum_{i=1}^{n_{\rm test}}L_{C(x_{i})}(a_{i};\alpha),$ where $L_{S}(a;\alpha)$ is defined in Lemma 2.1 and $a_{i}$ is the chosen action.
@@ Appendix B Proofs
Proof of Theorem 3.3.